To develop programs for business travelers staying at conventional hotels, a major hotel chain
commissioned a study of executives who play golf. The study revealed that 55% of the
executives admitted they have cheated at golf. Also, 20% of the executives admitted that they
had cheated at golf and had lied in business. Given an executive who had cheated at golf, what
is the probability the executive also had lied in business?
1
Expert's answer
2020-08-10T15:07:49-0400

Let A= executives admitted they have cheated at golf and

B= had lied in business


Given:

P(A)= 0.55

P(A and B)= 0.20


To find :

Given an executive who had cheated at golf, what is the probability the executive also had lied in business.

That is to find P(B/A).


Formula :

P(B/A)= "\frac{P(A and B)}{P(A)}"


solution:

P(B/A)= "\frac{0.2}{0.55}" = 0.3636 or 36.36%